please help solve I am stuck
2/5z + 8/z= 1

Answer by sudhanshu_kmr(1152)


2/5z + 8/z = 1
=> ( 2 + 5*8)/5z = 1
=> 42/5z = 1
=> 42 = 5z
=> z = 42/5
=> z = 8.4